### v1

# OpenSIR: Open-Ended Self-Improving Reasoner

Abstract
Recent advances in large language model (LLM) reasoning through reinforcement learning rely on annotated datasets for verifiable rewards, which may limit models’ ability to surpass human-level performance. While self-play offers a promising alternative, existing approaches depend on external verifiers or cannot learn open-endedly. We present Open-Ended Self-Improving Reasoner (OpenSIR), a self-play framework where an LLM learns to generate and solve novel problems by alternating teacher and student roles without external supervision. To generate novel problems, OpenSIR optimises for both difficulty and diversity, rewarding problems that challenge appropriately while exploring distinct concepts, enabling open-ended mathematical discovery. Starting from a single trivial seed problem, OpenSIR substantially improves instruction models: Llama-3.2-3B-Instruct advances from 73.9 to 78.3 on GSM8K, and from 28.8 to 34.4 on College Math, while Gemma-2-2B-Instruct rises from 38.5 to 58.7 on GSM8K. Our analyses reveal that OpenSIR achieves open-ended learning through co-evolving teacher-student roles that adaptively calibrate difficulty and drive diverse exploration, progressing autonomously from basic to advanced mathematics. The code is publicly available at https://github.com/EdinburghNLP/OpenSIR.
1 Introduction
Reinforcement learning with verifiable rewards (RLVR) drives recent advances in LLM reasoning. Recent works on DeepSeek-R1 (DeepSeek-AI et al., 2025) and OpenAI o1 (OpenAI, 2024) have shown that large-scale reinforcement learning improves reasoning capabilities. Yet these methods need extensive human-annotated data for reward signals, bottlenecking scalability and potentially limiting performance to human-level (Hughes et al., 2024b).
One promising direction to address these fundamental limitations is to generate synthetic training data through self-play, which demonstrated remarkable success in various games (Silver et al., 2016; 2017; Brown & Sandholm, 2019; FAIR et al., 2022), allowing systems to exceed human-level performance by learning from unambiguous reward signals (Silver et al., 2017; FAIR et al., 2022). Yet, mathematical reasoning poses a key challenge for self-play: unlike games that have clear rules and winners, generated mathematics problems lack the ground-truth answers to provide feedback signals. Recent works utilise external verifiers, such as compilers for coding tasks (Pourcel et al., 2024; Zhao et al., 2025) or game rules (Liu et al., 2025), while R-Zero (Huang et al., 2025) employs majority voting with basic repetition penalties. However, these approaches cannot achieve open-ended learning, the ability to continuously generate and pursue novel challenges without external supervision (Bauer et al., 2023; Hughes et al., 2024a), confining systems to known concepts instead of exploring diverse mathematical domains.
Figure 1: Overview of the OpenSIR framework. A single policy $\pi_{\theta}$ alternates between generating and solving novel problems without external supervision. Each training iteration consists of problem generation, solution sampling, scoring, and model update. Novelty is captured through both difficulty and diversity: problems must be challenging yet solvable, and they must explore new concepts. These dimensions together drive open-ended self-improvement in the LLM reasoning ability.
We present Open-ended Self-Improving Reasoner (OpenSIR), a method for training a policy $\pi_{\theta}$ to generate and solve novel problems without external supervision. OpenSIR uses self-play — a single policy $\pi_{\theta}$ alternates between teacher and student roles: the teacher generates problems, while the student solves them, with problem-solution pairs selected for reinforcement learning updates. We reward teachers for generating appropriately challenging problems for the students, using consistency and solution length across multiple solution attempts. OpenSIR achieves open-ended learning through embedding-based diversity rewards that drive continuous exploration of novel mathematical concepts.
[…]
Our qualitative analysis reveals OpenSIR succeeds through adaptive difficulty calibration and diversity-driven exploration. Problem difficulty is automatically calibrated throughout training, while the range of topics expands from basic to advanced mathematics (§4.1). Generating harder problems risks invalidity, requiring a balance between challenge and correctness (§4.2). Diversity rewards incentives generate problems spanning varied mathematical concepts (§4.3). Teacher-student co-evolution proves essential: without teacher training, models cannot generate appropriate challenges or explore new topics (§4.4).
2 Open-Ended Self-Improving Reasoner
Figure 1 illustrates the Open-Ended Self-Improving Reasoner (OpenSIR), a self-play framework in which a policy $\pi_{\theta}$ learns to both generate and solve novel mathematical problems without external supervision. We use reinforcement learning to optimise two roles within one policy: the teacher, which creates new problems, and the student, which solves them. This open-ended approach enables the policy to bootstrap its learning and discover new and diverse challenges without annotated data. Each training iteration involves four phases:
  1. 1.
    Problem generation2.1): The teacher proposes new problems by conditioning on reference problems from an accumulated pool of previously generated problems;
  2. 2.
    Solution sampling2.2): The student attempts multiple solutions per problem, with majority voting determining the reference answer and solve rate measuring reliability;
  3. 3.
    Scoring2.3): We compute novelty scores for the teacher’s generated problems and correctness scores for the student’s solutions; and
  4. 4.
    Model update2.4): We update the policy’s parameters with role-specific rewards using the problem-solution pairs selected by the novelty scores.
In OpenSIR, we define novelty along two dimensions that together drive continuous open-ended learning. First, problems must have an appropriate level of difficulty. It should be challenging enough to promote learning but solvable enough to provide reliable training signals. Second, problems must explore diverse concepts, preventing the model from repeating learning on familiar concepts. This two-dimensional view of novelty ensures the model continuously expands both the depth and breadth of its mathematical reasoning abilities.
2.1 Problem Generation
At each iteration $t$, the policy $\pi_{\theta}$ generates $k$ groups of $G$ problems each, denoted as $q_{1:G}$ within each group, for a total of $M=k\times G$ problems. To generate these problems, we sample $k$ reference problems from a pool $\mathcal{P}_{t-1}$ of accumulated problems from previous iterations, where each reference problem serves as a seed for generating $G$ new problems. Each generated problem must explicitly include the mathematical concepts required for its solution. Problems with invalid formats are filtered out, and valid problems proceed to the solution-sampling phase. We initialise the problem pool $\mathcal{P}_{0}$ with a single trivial problem (“What is 1+1?”).
2.2 Solution Sampling
Let $a_{j}$ denote the parsed answer from solution attempt $o_{j}$. We select the most common answer across attempts as the reference answer $a^{*}$. We then compute the solve rate for each problem to determine the reliability of the answers. For brevity, we denote $s_{q_{i}}=\text{SolveRate}(q_{i})$ when referring to the solve rate of problem $q_{i}$.
$\text{SolveRate}(q_{i})=\frac{\text{count}(a^{*})}{G}\quad\text{where}\quad a^{*}=\underset{a\in a_{1:G}}{\arg\max}\;\text{count}(a),$ (1)
In Eq.˜1, $\text{count}(a)$ denotes the number of times answer $a$ appears. The solve rate quantifies answer reliability. High solve rates indicate reliable reference answers due to solution convergence, while low solve rates suggest inconsistent solutions that may indicate flawed problem formulations.
2.3 Scoring
We evaluate the quality of generated problems and solutions with different scoring functions. The teacher’s problems are scored based on difficulty and diversity, while the student’s receive scores for correctness. Additionally, both roles incorporate format scores to ensure parseable outputs.
2.3.1 Teacher Scoring
We capture novelty through two fundamental dimensions: difficulty and diversity. We measure difficulty using solvability to ensure problems remain appropriately challenging and solution length to encourage multi-step reasoning, as these provide complementary signals about problem difficulty. Diversity is promoted through embedding distance, which encourages exploration of varied mathematical concepts. These components form a unified novelty score that guides problem generation.
Solvability ($\text{score}_{\text{sol}}$).
The solvability score identifies problems with appropriate challenge. We use solve rate as a proxy for solvability—problems with $s_{q_{i}}>s_{\max}$ are likely too easy, while those with $s_{q_{i}}<s_{\min}$ are either too difficult or malformed. We employ a triangular scoring function that peaks at the optimal solve rate and decreases linearly as problems become too easy or too hard.
We define the solve rate range as $s_{\min},s_{\max}]$. Easy problems ($s_{q_{i}}>s_{\max}$) fail to challenge the model, while problems that are too hard or malformed ($s_{q_{i}}<s_{\min}$) offer minimal training value.
Formally, for $s_{q_{i}}\in[0,1]$, let $s_{\text{mid}}=(s_{\min}+s_{\max})/2$ be the midpoint:
$\text{score}_{\text{sol}}(q_{i})=\begin{cases}1-\alpha|s_{q_{i}}-s_{\text{mid}}|&\text{if }s_{q_{i}}\in[s_{\min},s_{\max}],\\ 0&\text{otherwise}\end{cases}$ (2)
where $\alpha=\left(1-1/G\right)/(s_{\text{mid}}-s_{\min})$ is the slope coefficient, with $G$ being the number of solution attempts. The score peaks at the midpoint $s_{\text{mid}}$ and decreases to $1/n$ at the boundaries.
This creates a symmetric triangular score centred at the midpoint of the solve rate range, giving a maximum score for problems with moderate difficulty and progressively less score as the solve rate approaches either boundary.
Solution Length ($\text{score}_{\text{len}}$).
Solution length complements solvability by measuring problem complexity. Problems requiring multi-step reasoning typically elicit longer solutions. We score problems using the average length of student solutions:
$\text{score}_{\text{len}}(q_{i})=\min\left(\frac{\bar{l}(q_{i})}{l_{\text{base}}},\frac{l_{\text{cap}}}{l_{\text{base}}}\right)$ (3)
where $\bar{l}(q_{i})$ denotes average solution length for problem $q_{i}$, $l_{\text{base}}$ is a normalisation factor (defaults to 1000 tokens), and $l_{\text{cap}}$ prevents outliers from dominating the scoring signal. This score complements the solvability score (see Appendix C.1).
Diversity ($\text{score}_{\text{div}}$).
We compute the semantic distance between each new problem and the existing problem pool:
$\text{score}_{\text{div}}(q_{i})=\min_{q^{\prime}\in\mathcal{P}_{t-1}}d(e_{q_{i}},e_{q^{\prime}})$ (4)
where $e_{q_{i}}$ and $e_{q^{\prime}}$ represent problem embeddings obtained from a pre-trained encoder, and $d(\cdot,\cdot)$ denotes cosine distance. This score maximises when a problem is semantically distant from all existing problems in the pool.
Format ($\text{score}_{\text{fom}}^{T}$).
The format score ensures proper problem structure. Generated problems must be enclosed in <question> tags with concepts listed in <concepts> tags (maximum three concepts). We assign $\text{score}_{\text{fom}}^{T}(q_{i})=1$ for correct formatting and $\text{score}_{\text{fom}}^{T}(q_{i})=0$ otherwise.
Novelty Score.
We combine these components into a novelty score capturing both difficulty and diversity:
$\text{score}_{\text{novel}}(q_{i})=\alpha\text{score}_{\text{sol}}(q_{i})+\lambda\text{score}_{\text{len}}(q_{i})+\gamma\text{score}_{\text{div}}(q_{i})+\delta\text{score}_{\text{fom}}^{T}(q_{i})$ (5)
where $\alpha$, $\lambda$, $\gamma$, $\delta$ are hyperparameters that control the relative importance of each component. This novelty score is used to select high-quality problem-solution pairs for training.
2.3.2 Student Scoring
The student’s score is based on solution correctness. For each solution attempt, we evaluate correctness by comparing the parsed answer against the reference answer from majority voting.
Format ($\text{score}_{\text{fom}}^{S}$).
[…]
Correctness Score.
The student’s correctness score combines accuracy with the format score:
$\text{score}_{\text{correct}}(o_{j},a_{j})=\mathbf{1}[a_{j}=a^{*}]+\delta\text{score}_{\text{fom}}^{S}(o_{j})$ (6)
where $\mathbf{1}[a_{j}=a^{*}]$ is an indicator function that equals 1 when parsed answer $a_{j}$ from outcome $o_{j}$ matches the reference answer $a^{*}$, and 0 otherwise. This correctness score evaluates both solution accuracy and proper formatting.
2.4 Model Update
After computing novelty scores, we select $B$ high-quality samples from valid problems for reinforcement learning, allocating half to problem generation and half to solution solving. For teacher training, we choose problem groups with highest $\text{score}_{\text{novel}}$ variance to ensure diverse training signals. For student training, we select problems with the highest novelty scores to provide maximal training value.
We optimise the policy using $\pi_{\theta}$ with an objective similar to Group Relative Policy Optimization (GRPO) (Shao et al., 2024), adapted for on-policy training to ensure stability (Chen et al., 2025):
$\mathcal{J}(\theta)=\mathbb{E}_{\begin{subarray}{c}q_{1:G}\sim\pi_{\theta}(\cdot|p_{T})\\ o_{1:G}\sim\pi_{\theta}(\cdot|q_{i},p_{S})\end{subarray}}\left[\sum_{r\in\{T,S\}}\frac{1}{G}\sum_{i=1}^{G}A_{i}^{r}\right]-\beta\mathbb{D}_{KL}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)$ (7)
where $p_{T}$ and $p_{S}$ are the teacher and student prompts respectively, $r\in\{T,S\}$ refers to teacher and student, $\mathbb{D}_{KL}$ denotes the KL divergence, $\pi_{\mathrm{ref}}$ refers to the initial model before training. The advantage for each role $r\in\{T,S\}$ is computed as:
$A_{i}^{r}=\frac{R_{i}^{r}-\operatorname{mean}\left(R_{1:G}^{r}\right)}{\operatorname{std}\left(R_{1:G}^{r}\right)}.$ (8)
We define role-specific rewards $R_{i}^{T}$ and $R_{j}^{S}$ using the scoring functions from Section 2.3:
$R_{i}^{T}=\text{score}_{\text{novel}}(q_{i}),\quad R_{j}^{S}=\text{score}_{\text{correct}}(o_{j},a_{j})$ (9)
All valid problems are then added to the problem pool $\mathcal{P}_{t}$ for future iterations.
3 Experiments
3.1 Training Setup
We experiment with three instruction-tuned models: Llama-3.2-3B-Instruct (Dubey et al., 2024), Gemma-2-2B-Instruct (Team et al., 2024), and Qwen-2.5-3B-Instruct (Team, 2024) with GRPO (Shao et al., 2024). We use a learning rate of $3\times 10^{-7}$ and 10 warm-up steps. The KL divergence coefficient is set to $10^{-4}$ and the batch size is 256. To compare models trained on the same number of problem-solution pairs, we train the GRPO baselines with 100 steps, and OpenSIR for 200 steps since OpenSIR allocates half of its training budget to problem generation. Clipping is not applied since we strictly use on-policy samples. Each experiment is run with three random seeds. We provide full training details in Appendix D.1.
3.2 Dataset and Evaluation Setup
We evaluate method on five mathematical benchmarks: GSM8K (Cobbe et al., 2021), MATH-500 (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), OlympiadBench (He et al., 2024), and College Math (Tang et al., 2024).
We use sampling temperature $0.6$ and top-p $0.95$. The maximum response length is set to 4,096 tokens. We report the average performance over 16 generations (avg@16). Answer extraction and comparison are performed using the math_verify library.
3.3 Baselines
(1) Base We evaluate the instruction-tuned models using zero-shot prompting, where models generate step-by-step reasoning and provide final answers without additional training.
[…]
3.4 Main Results
Model GSM8K MATH-500 Minerva College OlympiadBench Avg.
Math
Llama-3.2-3B-Instruct
Base 73.94 42.86 15.21 28.78 13.09 34.78
[…]
OpenSIR 78.28 46.22 17.46 34.42 15.72 38.42+3.6
Gemma-2-2B-Instruct
Base 38.50 16.51 10.09 19.11 3.00 17.44
[…]
OpenSIR 58.03 24.75 9.51 23.36 3.15 23.76+6.3
Qwen-2.5-3B-Instruct
Base 84.43 65.36 25.23 48.22 27.94 50.24
[…]
OpenSIR 85.38 65.87 25.96 48.74 28.33 50.85+0.6
Table 1: The avg@16 performance on five mathematical benchmarks. OpenSIR consistently outperforms GRPO baselines across model architectures despite generating training data through self-play from a single seed problem, while GRPO baselines use over 7,000 human-annotated examples.
Table 1 demonstrates that OpenSIR achieves substantial gains over the base instruction models, improving Llama-3.2-3B-Instruct by 3.6 points and Gemma-2-2B-Instruct by 5.9 points on average accuracy. While OpenSIR demonstrates huge improvements with Llama-3.2-3B-Instruct and Gemma-2-9b-Instruct, Qwen-2.5-3B-Instruct demonstrates a relatively minimal gains (+0.3). Notably, GRPO baselines exhibit similarly marginal improvements with this model, indicating the effect is not specific to OpenSIR. The limited improvement aligns with observations of potential benchmark contamination (Wu et al., 2025).
OpenSIR outperforms all GRPO baselines without using human-annotated training data. GRPO baselines require over 7,000 labeled examples, yet OpenSIR generates its own training problems through self-play, starting from a single trivial seed problem. As we reveal later in our analysis, the success of OpenSIR can be attributed to its ability to explore diverse mathematical concepts, and calibrating difficulty adaptively to maintain optimal challenge levels (§4.1). These capabilities enable OpenSIR to self-improve and expand its skills without external training data, achieving open-ended learning.
4 Ablations and Analyses
We perform a series of ablation studies and qualitative analyses on Llama-3.2-3B-Instruct to dissect the contribution of each key component in the OpenSIR framework. Our analysis investigates: (1) the evolution of problem difficulty and diversity over training (§4.1), (2) the effect of solve rate thresholds on the difficulty-validity trade-off (§4.2), (3) the impact of diversity rewards on promoting exploration of novel problem types (§4.3), and (4) the necessity of dual-role training (§4.4).
4.1 Evolution of Problem Difficulty and Diversity
Figure 2: Evolution of problem difficulty, validity, and topic diversity during OpenSIR training. (Left) Human evaluation results showing difficulty rankings (1-5 scale where 1=easiest, 5=hardest) and number of invalid problems for GSM8K, MATH, and problems generated at steps 0, 100, and 200 of training. Invalid problems are those with logical flaws, missing information, or ambiguities. (Right) Distribution of mathematical topics across training stages, demonstrating the increasing diversity of generated problems from step 0 to step 200.
We track how difficulty and diversity evolve during training through human evaluation. We sample 20 problems from three OpenSIR training checkpoints (steps 0, 100, 200) and 20 each from GSM8K and MATH. Annotators evaluate mixed sets of five problems (one per source), identifying topics, assessing validity, and ranking difficulty. Figure 2 shows average difficulty rankings (1=easiest, 5=hardest); see Appendix B for full annotation instructions.
Figure 2 (left) reveals a V-shaped difficulty trend across training stages. Problems start at 3.4 difficulty, drop to 3.0 at midpoint, then rise to 3.8. This pattern reflects OpenSIR’s self-calibration: the model first generates overly difficulty problems, then learns appropriate difficulty, and finally increases challenge as its solving capabilities improve. The model also generates increasingly valid problems during training — validity improves from below 50% initially to 95% (19 of 20 problems) by the end.
Figure 2 (right) shows topic diversity expansion across training. OpenSIR progresses from basic topics (algebra, arithmetic, geometry) to advanced domains including calculus and optimisation, eventually incorporating trigonometry, statistics, and other mathematical areas. This progression demonstrates OpenSIR’s capacity for autonomous exploration of diverse mathematical concepts. Appendix A.2 provides detailed case studies that illustrate this evolution.
4.2 Difficulty-Validity Trade-off
Model Acc Validity Solve Rate
OpenSIR0.5 38.42 70.82 89.82
OpenSIR0.3 36.81 52.32 81.38
OpenSIR0.1 35.97 42.31 78.31
Table 2: Performance, problem validity, and solve rate across different lower solve-rate thresholds, with the upper threshold fixed at 0.9 for all variants. Validity and solve rate are estimated using GPT-5. Lower thresholds produce harder problems but significantly more invalid ones, ultimately reducing overall performance.
We investigate the difficulty-validity trade-off by training OpenSIR variants with lower solve-rate thresholds of 0.1, 0.3, and 0.5, keeping the upper threshold at 0.9 From each variant, we sample 300 problems and assess quality with GPT-5 (OpenAI, 2025a) using 8 responses per problem. We measure validity by comparing GPT-5’s majority answer to our reference answer and difficulty by GPT-5’s solve rate.
Table 2 reveals a clear trade-off between validity and difficulty. While lowering the threshold from 0.5 to 0.1 produces moderately harder problems (GPT-5 solve rate decreases from 89.82% to 78.31%), validity plummets from 70.82% to 42.31%. This suggests that problems with very low solve rates frequently contain errors rather than representing genuine mathematical challenges. performance consistently drops with lower thresholds, supporting our selection of 0.5 as the lower threshold for the solvability reward.
Besides solve-rate thresholds, we find that rewarding longer solutions provides another mechanism for promoting problem complexity that encourage sophisticated multi-step problems (Appendix C.1).
4.3 Impact of Diversity Rewards
Figure 3: t-SNE visualization of problem embeddings showing the effect of diversity reward on problem distribution. With diversity reward, problems explore broader regions of the embedding space compared to the clustered distribution without diversity reward.
Model Acc # Concepts
w diversity 38.42 5914
w/o diversity 36.45 3328
Table 3: OpenSIR performance with and without diversity reward. Exploring diverse mathematical concepts through the diversity reward improves both accuracy and concept coverage, showing that variety in problem types is crucial for self-improvement.
We analyse the impact of the diversity reward on problem diversity through problem embeddings, n-gram similarity, and concept overlap. Figure 3 visualises the problem embeddings with t-SNE, where red points represent problems without diversity reward, cyan points show problems with diversity reward, gold indicates MATH dataset problems, and purple marks GSM8K dataset problems. Without diversity rewards, problems cluster in narrow regions, generating similar types repeatedly and failing to achieve open-ended exploration. With diversity rewards, problems spread across the embedding space, reaching areas beyond MATH and GSM8K training sets. Further analysis of n-gram similarity and concept overlap support these findings, demonstrating consistent patterns of greater dispersion and novelty (Appendix A.3).
Table 3 empirically confirms the importance of diversity rewards, showing that removing diversity rewards reduces average performance by 1.97 (from 38.42 to 36.45). It also shows that the number of unique concepts has dropped significantly (from 5914 to 3328). This demonstrates that without diversity rewards, the model generates repetitive problems with limited learning value, constraining the teacher’s ability to present varied mathematical challenges to the student. Incorporating diversity rewards thus enables exploration of novel problems beyond existing datasets, supporting open-ended learning where the model continuously discovers new challenges rather than repeating known concepts. Notably, this improvement is robust to the choice of diversity metric (Appendix C.2), with different measurement approaches yielding comparable results.
4.4 Importance of Dual-Role Training
Trained Acc Avg.
Roles Solve Rate
Both 38.42 72.20 ($\pm$4.49)
Student 35.89 64.56 ($\pm$17.37)
Table 4: Accuracy and average solve rate with standard deviation ($\pm$) for OpenSIR with teacher training (Both) versus without teacher training (Student only). Joint training achieves higher accuracy and remarkably stable problem difficulty (much lower solve rate variance), demonstrating that teacher training enables calibrated problem generation at optimal difficulty levels for effective learning.
We evaluate the contribution of the joint teacher-student training by testing a variant where only the student is updated while the teacher remains fixed at its initial state. Table 4 shows that accuracy drops significantly from 38.42 to 35.89 when only the student is trained. This demonstrates that effective self-play requires both components to co-evolve.
Without teacher training, generated problems become harder (solve rate drops from 72.20 to 64.56) and drift from the optimal 70% target solve rate established in Section 4.2. More critically, solve rate variance increases tremendously (from $\pm$4.49 to $\pm$17.37), indicating highly inconsistent difficulty during training. This poorly calibrated curriculum explains the performance drop: the fixed teacher cannot adapt to the student’s evolving capabilities, whereas joint training enables continuous difficulty calibration at the optimal challenge level.
5 Related Work
Self-play.
Self-play achieved superhuman performance in games without human data, from AlphaGo (Silver et al., 2016; 2017), StarCraft II (Vinyals et al., 2019), Poker (Brown & Sandholm, 2019), DotA (OpenAI et al., 2019), and Diplomacy (FAIR et al., 2022). Baker et al. (2019) show that agents can discover complex strategies with self-play, suggesting it is a promising avenue for continuous open-ended learning. Recent works apply self-play to LLM reasoning: Absolute Zero (Zhao et al., 2025) and Spiral (Liu et al., 2025) rely on external verifiers or game rules that limit their use beyond specific domains. R-zero (Huang et al., 2025) attempts verifier-free self-play but uses only repetition penalties without a mechanism to encourage exploration, constraining open-ended learning. In contrast, OpenSIR generates and solves problems without external supervision while actively promoting diversity to enable continuous discovery of novel mathematical concepts.
Reinforcement Learning with Verifiable Feedback (RLVF).
RLVF drives recent advances in LLM reasoning (OpenAI, 2024; 2025b; DeepSeek-AI et al., 2025) but requires extensive human-annotated data for verifiable reward signals (Zeng et al., 2025), creating scalability bottleneck and potentially limiting performance to human-level. Recent works show that moderate-difficulty training samples provide optimal learning signals (Zheng et al., 2025; Sun et al., 2025), while diverse problem types enhance mathematical reasoning (Akter et al., 2025; Chen et al., 2025). These insights directly motivate OpenSIR to optimise for appropriate difficulty calibration and diversity-driven exploration, enable models to learn math reasoning open-endedly without human supervision.
6 Conclusions
We present OpenSIR, a self-play framework that enables LLMs to autonomously learn to generate and solve novel problems without external supervision. Starting from only a single trivial math problem, our framework outperforms GRPO-trained models that utilise thousands of human annotations across diverse model families. This approach demonstrates that models can effectively bootstrap mathematical reasoning through recursive self-improvement, eliminating dependence on extensive curated datasets. Our analysis reveals that OpenSIR succeeds by combining difficulty calibration and diversity rewards to create an adaptive curriculum where models continuously discover and master increasingly challenging mathematical concepts. Overall, OpenSIR represents a compelling paradigm for open-ended autonomous mathematical reasoning development, enabling models to recursively expand their capabilities beyond the boundaries of human-annotated data.
Acknowledgments
The authors would like to thank Aryo Pradipta Gema, Neel Rajani, Rohit Saxena (in alphabetical order) for the helpful discussions and feedback on the manuscript.
Appendix A Extended Results and Analysis
A.1 Full Results
Models Seed GSM8K MATH-500 Minerva College Olympiad- Avg.
Math Bench
Llama-3.2-3B-Instruct
Base - 73.94 42.86 15.21 28.78 13.09 34.78
[…]
43 79.62 44.56 16.64 33.35 14.52 37.74
44 79.93 45.91 15.83 33.32 14.46 37.89
[…]
43 76.51 45.23 15.95 32.87 13.85 36.88
44 75.93 45.52 15.95 32.95 14.23 36.92
[…]
OpenSIR 42 77.82 46.38 17.72 34.24 15.46 38.32
43 78.58 45.91 17.23 34.58 15.86 38.43
44 78.43 46.38 17.44 34.45 15.84 38.51
[…]
Gemma-2-2B-Instruct
Base - 38.50 16.51 10.09 19.11 3.00 17.44
[…]
43 58.86 19.21 7.96 20.77 3.08 21.98
44 59.06 19.36 7.76 20.42 3.37 21.99
[…]
43 53.94 22.53 7.90 15.08 3.11 20.51
44 59.01 23.44 8.02 18.15 3.57 22.44
[…]
OpenSIR 42 58.68 24.09 8.89 22.29 2.99 23.39
43 58.36 25.69 10.73 26.14 3.23 24.83
44 57.03 24.49 8.89 21.66 3.24 23.06
[…]
Qwen-2.5-3B-Instruct
Base - 84.43 65.36 25.23 48.22 27.94 50.24
[…]
43 85.16 65.80 24.84 48.46 28.50 50.55
44 84.96 66.10 24.75 48.40 28.23 50.49
[…]
43 84.19 65.64 25.14 48.20 27.98 50.23
44 84.49 66.30 24.59 48.29 28.57 50.45
[…]
OpenSIR 42 85.43 66.17 26.49 48.88 28.86 51.17
43 85.26 65.64 25.30 48.62 28.30 50.62
44 85.44 65.79 26.08 48.72 27.83 50.77
[…]
Table 5: Math reasoning evaluation results with individual seed reporting. We report avg@16 per problem for each seed (s-42, s-43, s-44) and their average with standard deviation as superscript.
We provide the full results of all seeds in Table 5.
A.2 Case Study
This section provides further analysis of question-solution pairs during training.
As discussed in Section 4.1, the model generates predominantly invalid problems early in training. Majority of these problems, primarily involve simple mathematical concepts like arithmetic, fail due to missing information (Figures 4 and 5). When attempting complex topics like optimisation, which are rare in the beginning, the model produces problems with missing information and fundamental formulation errors (Figure 6). This reveal the model has limited understanding of underlying mathematical concepts. Invalid problems tend to exhibit low solve rates ($\leq$ 0.25) and correspondingly receive lower rewards, helping the model learn to generate valid problems. Consequently, invalid problems decrease rapidly across training (§4.1).
However, not all problems with low solve rates are invalid (§4.2). We find that some problems involving certain topics that are challenging for the model, such as geometric series, persistently exhibit low solve rates (Figures 8. The model struggles with exponentiation calculations, resulting in poor performance on geometric series problems. This reveals a fundamental trade-off in OpenSIR: while higher solve rate thresholds effectively filter out invalid problems, they inevitably discourage exploration of genuinely difficult topics. Since these problems have low solvability scores, they are likely to not receive sufficient encouragement to further explore these topics.
In later training stages, we observe OpenSIR gradually expanding into advanced mathematical domains. After 100 training steps, the model starts to generate problems involving concepts like optimisation (Figure 9), calculus (Figure 10), trigonometry-based physics (Figure 11), probability (Figures 12), among others. While these advanced problems yield lower solve rates, which indicate the model has a limited understanding of these domains, they achieve high novelty scores with large semantic distances and longer solutions. This progression validates how novelty rewards in OpenSIR drive exploration of diverse mathematical concepts, enabling open-ended learning.
A concert venue charges an admission price per seat and also offers a VIP ticket package that includes food, drinks, and other perks. If the food and other perks are included in the VIP ticket price and are worth \$10 per person, and a group of friends want to buy the same number of VIP tickets as they would with regular tickets, what is the minimum admission price per regular ticket to make the total cost of the VIP tickets equal to or less than the total cost of the regular tickets?
Figure 4: An invalid arithmetic question generated in step 0 with solve rate of 0.25. This question is invalid since the VIP tick price is not provided, and therefore, it’s impossible to calculate the minimum regular ticket price.
Find the percentage difference in the cumulative growth of two continuously compounded interest functions after 5 years: $A=P*e^{(rt)}$, where A is the amount of money accumulated after n years, including interest, P is the principal amount, r is the annual interest rate, and t is the time the money is invested for.
Figure 5: An invalid arithmetic question generated in step 0 with solve rate of 0.125. This question is invalid since the two interest rates and principal amounts are not provided. Hence, it’s impossible to calculate the percentage difference with just the general formula provided.
Consider two positive integers m and n (m $\geq$ n). Given a 2x2 matrix of numbers where each element is a non-negative integer, find the maximum value of the following expression: $ax^{2}+by^{2}+cxy$, where a, b, and c are constants, subject to the constraint that the sum of any two elements in each row and column of the matrix are equal. What is the maximum possible value of $ax^{2}+by^{2}+cxy$?
Figure 6: An invalid optimisation question generated in step 0 with solve rate of 0.125. This question is invalid because there are missing information about the constants a, b, and c. There are also ambiguities in the question, such as the role of m and n in the problem. It also did not explain what the elements of the matrix represent. Lastly, it contains problem formulation errors, specifically failing to specify constraints that ensure bounded solutions, demonstrating insufficient understanding of optimization problem structure.
Consider two positive integers m and n (m $\geq$ n). Given a 2x2 matrix of numbers where each element is a non-negative integer, find the maximum value of the following expression: $ax^{2}+by^{2}+cxy$, where a, b, and c are constants, subject to the constraint that the sum of any two elements in each row and column of the matrix are equal. What is the maximum possible value of $ax^{2}+by^{2}+cxy$?
Figure 7: An invalid optimisation question generated in step 0 with solve rate of 0.125. This question is invalid because there are missing information about the constants a, b, and c. There are also ambiguities in the question, such as the role of m and n in the problem. It also did not explain what the elements of the matrix represent. Lastly, it contains problem formulation errors, specifically failing to specify constraints that ensure bounded solutions, demonstrating insufficient understanding of optimization problem structure.
Question A company is adding new employees to its workforce. Initially, the company has 500 employees and adds 10 new employees every quarter. If this trend continues for the next 5 years, and each quarter the company adds a fixed percentage increase in the number of new employees (starting with 10%), how many new employees can the company be expected to add during the first 2 years?
Solution (Incorrect) To find the total number of new employees added in the first 2 years, we need to calculate the number of new employees added each quarter for the first 2 years and then sum them up. First, let’s calculate the number of quarters in 2 years: 2 years * 4 quarters/year = 8 quarters. The company starts with 10 new employees and adds 10% more each quarter. To find the total number of new employees added each quarter, we can use the formula for compound interest: $A=P*(1+r)^{n}$ where: A = total number of new employees after n quarters P = initial number of new employees (10) r = percentage increase per quarter (10% or 0.1) n = number of quarters However, since we want to find the total number of new employees added each quarter, we need to find the difference between A and the initial number of new employees (10). We can rewrite the formula as: $A-P=P*(1+r)^{n}$ $A-10=10*(1+0.1)^{n}$ $A=10*(1.1)^{n}$ Now, we can calculate A for each quarter: Quarter 1 (n = 1): A = $10*(1.1)^{1}=11$ Quarter 2 (n = 2): A = $10*(1.1)^{2}=12.1$ Quarter 3 (n = 3): A = $10*(1.1)^{3}=13.21$ Quarter 4 (n = 4): A = $10*(1.1)^{4}=14.41$ Quarter 5 (n = 5): A = $10*(1.1)^{5}=15.71$ Quarter 6 (n = 6): A = $10*(1.1)^{6}=17.31$ Quarter 7 (n = 7): A = $10*(1.1)^{7}=19.33$ Quarter 8 (n = 8): A = $10*(1.1)^{8}=21.59$ Now, we can calculate the total number of new employees added in the first 2 years by summing up the number of new employees added each quarter: Total = 11 + 12.1 + 13.21 + 14.41 + 15.71 + 17.31 + 19.33 + 21.59 Total = 119.66 Since we can’t have a fraction of an employee, we can round down to the nearest whole number. Therefore, the final answer is: $\boxed{119}$
Figure 8: A valid arithmetic problem involving geometric series with a solve rate of 0.125. The solution requires either summing quarterly employee additions or applying the geometric series formula: $a\cdot\frac{r^{n}-1}{r-1}$. Llama-3.2-3B-Instruct struggles with exponentiation calculations, resulting in poor performance on geometric series problems.
A farmer has 10 different types of vegetables to store in a warehouse with 1,000 square meters of floor space. There are four container sizes available: - Small (5 sq m): Maximum 50 available - Medium (10 sq m): Maximum 40 available - Large (15 sq m): Maximum 30 available - Extra-large (20 sq m): Maximum 25 available The vegetables have different storage requirements: - 3 bulky vegetables (pumpkins, watermelons, cabbages) require containers of at least 15 sq m - 4 medium vegetables (tomatoes, peppers, eggplants, zucchini) require containers of at least 10 sq m - 3 small vegetables (carrots, onions, potatoes) can fit in any container size Each vegetable type must be stored in at least one container. What is the maximum number of containers that can be used while satisfying all constraints and not exceeding 1,000 sq m total space?
Figure 9: A valid optimisation problem with a solve rate of 0.375 generated at step 124.
Find the equation of the curve y = f(x) where the derivative is given by f(x) = ($3x^{2}-x-2$)/2x and the curve passes through the point (2, 3).
Figure 10: A valid calculus problem with a solve rate of 0.375 generated at step 156.
A golfer hits a ball from the top of a 50-meter high cliff with an initial velocity of 30 m/s at an angle of 45 degrees above the horizontal. What is the horizontal distance traveled by the ball when it hits the ground?
Figure 11: A valid physics problem that involves trigonometry with a solve rate of 0.5 generated at step 172.
Consider a randomly ordered sequence of $n=3q$ distinct integers $\{a_{1},a_{2},\ldots,a_{3q}\}$ where $q$ is a positive integer. Define $f$ as the number of adjacent pairs $(a_{i},a_{i+1})$ in the sequence where both integers have the same remainder when divided by 3 (i.e., $a_{i}\bmod 3=a_{i+1}\bmod 3$). If the integers 11 through $3q$ are randomly permuted to form this sequence, what is the expected value of $f$?
Figure 12: A valid probability problem with a solve rate of 0.25 generated at step 188.
A.3 Further Analysis on Questions Diversity
Figure 13: Heatmap visualisation of n-gram similarity (ROUGE-L scores) and concept overlap between generated problems at training steps 0, 100, 200 and reference datasets (MATH, GSM8K). Top row: with diversity reward; Bottom row: without diversity reward. With diversity reward incorporated, the generated problems exhibit low textual similarity and minimal concept overlap, demonstrating effective exploration of diverse problem types.
Figure 13 presents n-gram similarity and concept analysis. We compute ROUGE-L scores between problem texts and extract mathematical concepts using GPT-5 from problems at steps 0, 100, and 200, as well as from the MATH and GSM8K training sets. With diversity rewards (top row), problems maintain low ROUGE-L scores and minimal concept overlap both across training stages and with MATH/GSM8K. Without diversity rewards (bottom row), both textual similarity and concept overlap increase, confirming limited exploration of new problem types.
Appendix B Annotation Details
One of the authors prepare the samples for annotation, and the rest of the authors annotated the samples with the instructions provide in Figure 14.
You will be presented with multiple sets of 5 math problems to evaluate. For each set, please complete the following three-step annotation process. # Step 1: Identify Topics For each problem, identify ALL relevant mathematical topics from the following list: - Algebra - Geometry - Calculus - Probability - Statistics - Number Theory - Combinatorics - Optimization - Arithmetic - Discrete Math - Trigonometry # Step 2: Assess Validity For each problem, determine if it is valid or invalid: - Valid: The problem is logically sound, clearly stated, and can be answered with the given information - Invalid: The problem contains logical flaws, contradictions, insufficient information, or ambiguities that prevent a proper solution # Step 3: Rank Difficulty Rank all 5 problems from easiest to hardest. Provide your ranking as a sequence of problem numbers. Example: [3, 1, 5, 2, 4] means problem 3 is the easiest and 4 is the hardest. Consider these factors when assessing difficulty: - Number of steps required - Complexity of concepts involved - Level of mathematical knowledge needed - Computational complexity # Response Format Provide your annotations as a JSON list where each element represents one problem set. Here are some examples:
    "set_id": "SET_1",
    "problems": {
      "1": {"topics": ["Algebra", "Calculus"], "valid": true},
      "2": {"topics": ["Geometry"], "valid": false},
      "3": {"topics": ["Probability"], "valid": true},
      "4": {"topics": ["Number Theory"], "valid": true},
      "5": {"topics": ["Arithmetic"], "valid": true}
    "difficulty_ranking": [5, 3, 1, 2, 4]
    "set_id": "SET_2",
    "problems": {
      "1": {"topics": ["Statistics"], "valid": true},
      "2": {"topics": ["Discrete Math"], "valid": true},
      "3": {"topics": ["Optimization"], "valid": true},
      "4": {"topics": ["Algebra"], "valid": false},
      "5": {"topics": ["Geometry", "Algebra"], "valid": true}
    "difficulty_ranking": [1, 2, 5, 3, 4]
Figure 14: The instruction provided to the annotators to annotate problems.
Appendix C Additional Ablations
C.1 Solution Length Reward Increases Problem Complexity
Model Question Solution Acc
Length Length
w/ length 207 387 38.42
w/o length 150 238 37.86
Table 6: Comparison of OpenSIR performance with and without solution length reward. Solution length reward improves OpenSIR accuracy and increases average question and solution lengths.
We investigate the impact of the solution length reward in OpenSIR. Table 6 shows this reward improves performance from 37.86% to 38.42%. It also increases the average question length (from 150 to 207 tokens) and solution lengths (from 238 to 387 tokens). By examining the generated questions manually, we find that the policy tends to generate more sophisticated problems involving advanced concepts with this reward, such as linear programming and optimization, which naturally require longer multi-step solutions to solve. These results demonstrate that the solution length reward effectively guides the policy toward generating more complex problems, which in turn leads to better performance.
C.2 Robustness to Diversity Measurements
Reward Acc # Concepts
Embedding 38.42 5914
Concepts 38.26 6213
Table 7: Comparison of diversity measurement approaches in OpenSIR. Despite slight differences in concept coverage, both embedding-based and concept-based diversity rewards yield nearly identical accuracy, demonstrating the framework’s robustness to the choice of diversity metric.
We have established the necessity of diversity rewards in Section 4.3. In this section, we further investigate OpenSIR’s robustness to different diversity measurement approaches. We implement concept-based diversity by measuring diversity through the mathematical concepts of the problems (Lu et al., 2023; Havrilla et al., 2025). Formally, we define the concept diversity reward as:
$r_{\text{con}}(q)=\frac{|\mathcal{C}_{q}|-|\mathcal{C}_{q}\cap\mathcal{C}_{\mathcal{P}_{t-1}}|}{3}$ (10)
where $\mathcal{C}_{q}$ are the concepts in problem $q$ and $\mathcal{C}_{\mathcal{P}_{t-1}}=\bigcup_{q^{\prime}\in\mathcal{P}_{t-1}}\mathcal{C}_{q^{\prime}}$ represents the union of concepts from all problems in the existing pool. Since each problem contains at most three concepts, this reward calculates the fraction of new concepts introduced.
Table 7 shows that both embedding-based and concept-based diversity rewards achieve similar accuracy (38.42 vs 38.26), demonstrating the framework’s robustness to the choice of diversity metric. Beyond accuracy, we examine concept coverage, which refers to the number of unique mathematical concepts discovered during training, as a direct measure of exploratory diversity. As expected, concept-based diversity achieves slightly higher coverage (6,213 concepts) since it explicitly optimises for novel concept discovery. Surprisingly, embedding-based diversity attains comparable coverage (5,914 concepts), 95% of the concept-based approach, despite not tracking concepts explicitly. This suggests that maximising representational spread in embedding space effectively promotes novelty discovery, achieving open-ended learning.
Appendix D Implementation Details
D.1 Training Details
Category Hyperparameter Value
Trainer Learning rate $3\times 10^{-7}$
Optimiser AdamW (Loshchilov & Hutter, 2018)
Warmup steps 20
Training steps 100/200
KL loss coefficient $1\times 10^{-4}$
Gradient norm clipping 0.5
Seeds 42/43/44
GPUs 3 H100
[…]
Max prompt length 1024
Max solution length 2048
Number of rollouts per prompt 8
Temperature 1.0
Teacher Rewards Solvability weight ($\alpha$) 1.0
Solution length weight ($\lambda$) 1.0
Diversity weight ($\gamma$) 1.0
Format weight ($\delta$) 0.1
Embedding model Linq-Embed-Mistral (7B)
Student Rewards Accuracy weight 1.0
Format weight ($\delta$) 0.1
  • $\dagger$
    The number of rollouts seen for one gradient update.
Table 8: The training configurations for the experiments.
We implement OpenSIR based on the TRL framework (von Werra et al., 2020). Table 8 provides a summary of the training hyperparameters used in our experiments.
D.2 Prompts
We detailed the prompt for generating problems in Figure 15 and solving problems in Figure 16.
You are given a math problem: {Problem} Your task is to create a math problem that is conceptually different from the provided problem. The new problem must be answerable with a numerical value or mathematical expression. First, explain how your new problem differs conceptually from the original problem inside the <think>…</think> tags. Then, present your new problem inside the <problem>…</problem> tags. Finally, identify at most three math concepts required to solve your problem. Provide these concepts in a comma separated list inside the <concepts>…</concepts> tags.
Figure 15: Prompt for generating math problems. {Problem} is a placeholder for the reference problem sampled from the problem pool.
[…]
Figure 16: Prompt for generating solutions to math problems. {Problem} is a placeholder for the actual problem.